# LongLLaDA: Unlocking Long Context Capabilities in Diffusion LLMs

### v2

@@ 6 Related Work
Recently, Large Language Diffusion Models, or diffusion LLMs, have become a widely discussed topic in NLP research. After the theoretical simplification (Sahoo et al., 2024; Ou et al., 2024; Shi et al., 2024) and empirical verification (He et al., 2022; Gong et al., 2024), researchers scale the size of diffusion LLMs to billions of parameters (Nie et al., 2024; 2025; Ye et al., 2025) and demonstrate that diffusion LLMs can achieve comparable results with more promising performance in the reversal curse (Berglund et al., 2023). These immediately attract the attention of many more researchers. Significant research efforts have focused on adapting diffusion LLMs for multimodality, such as MMaDA (Yang et al., 2025), LLaDA-V (You et al., 2025), and LaViDa (Li et al., 2025), applying them to reasoning tasks, such as d1 (Zhao et al., 2025), DCoLTHuang et al. (2025), and LLaDA-1.5 (Zhu et al., 2025), and optimizing their efficiency (Ma et al., 2025; Hu et al., 2025; Wu et al., 2025), including dKV-Cache (Ma et al., 2025), Dimple (Yu et al., 2025), dLLM-Cache (Liu et al., ), FreeCache (Hu et al., 2025), Fast-dLLM (Wu et al., 2025), and so on. However, there is still no discussion on the long-context capability of diffusion LLMs.
@@
Length extrapolation, or length generalization, or context extension, is an important issue for LLMs (Press et al., 2022; Liu et al., 2025). The mainstream extrapolation research mainly focuses on adjusting position embedding, especially the widely used RoPE (Su et al., 2021). For example, Linear PI (Chen et al., 2023) first achieves LLMs’ length extrapolation by scaling position indices to the pre-training range with little fine-tuning. The NTK method (bloc97, 2023b; a; Peng et al., 2023) then scales the rotary base in RoPE (Su et al., 2021) to achieve plug-and-play length extrapolation. Subsequently, amplifying the rotary base and training on longer lengths has become the dominant approach for length extrapolation (Rozière et al., 2023; Xiong et al., 2023; Liu et al., 2023b; Ding et al., 2024). In addition, ReRoPE (Su, 2023), ReAttention (Liu et al., 2024b), and DCA (An et al., 2024a; b) also achieve plug-and-play extrapolation by limiting the relative position. In this paper, we still focus on the length extrapolation via NTK scaling (bloc97, 2023b; Liu et al., 2023b) in the inference stage, and try to reveal and explain the similarities and differences in length extrapolation between diffusion-based and auto-regressive LLM.
@@ Appendix A Preliminary: RoPE Extrapolation in Auto-regressive LLM
Rotary Position Embedding (RoPE) (Su et al., 2021) employs trigonometric functions to encode absolute positions in Q state $\bm{q}_{t}=\left[q_{t}^{(0)},\cdots,q_{t}^{(d-1)}\right]$ and K state $\bm{k}_{s}=\left[k_{s}^{(0)},\cdots,k_{s}^{(d-1)}\right]$. By leveraging the properties of rotation matrices, RoPE encodes relative position in the attention matrix $\bm{A}$, as shown in Equantion 2 and demonstrates superior performance, thus being widely adopted by many auto-regressive LLMs (Sun et al., 2024; Dubey et al., 2024; Yang et al., 2024).
$\bm{A}_{t,s}&=\left(\bm{q}_{t}\bm{R}_{t}\right)\left(\bm{k}_{s}\bm{R}_{s}\right)^{\top}=\bm{q}_{t}\bm{R}_{t-s}\bm{k}_{s}^{\top}\\ &=\sum\limits_{n=0}^{d/2-1}{\begin{bmatrix}q_{t}^{(2n)}\\ q_{t}^{(2n+1)}\end{bmatrix}^{\top}\begin{bmatrix}\cos{\theta_{n}(t-s)}&-\sin{\theta_{n}(t-s)}\\ \sin{\theta_{n}(t-s)}&\cos{\theta_{n}(t-s)}\end{bmatrix}\begin{bmatrix}k_{s}^{(2n)}\\ k_{s}^{(2n+1)}\end{bmatrix}}\\ &=\sum\limits_{n=0}^{d/2-1}{\begin{gathered}\left(q_{t}^{(2n)}k_{s}^{(2n)}+q_{t}^{(2n+1)}k_{s}^{(2n+1)}\right)\cos{\theta_{n}(t-s)}\\ -\left(q_{t}^{(2n)}k_{s}^{(2n+1)}-q_{t}^{(2n+1)}k_{s}^{(2n)}\right)\sin{\theta_{n}(t-s)}\end{gathered}}\end{aligned}\text{.}$ (2)
However, RoPE still faces the length extrapolation issue (Press et al., 2022). When RoPE-based auto-regressive LLMs are tested beyond the pre-trained context length, the perplexity rises significantly, and downstream performance drops sharply. The underlying causes and corresponding solutions can be attributed to two key properties of trigonometric functions: periodicity and monotonicity.
@@
Therefore, dimensions beyond $d_{\text{extra}}$ will encounter OOD position embedding when processing longer inputs and larger position indices in inference, leading to extrapolation issues (Liu et al., 2023b).
To enable LLM to handle unseen position indices, NTK methods (bloc97, 2023b; Xiong et al., 2023) scale the rotary base by a factor $\lambda$, reducing the rotary angle to achieve position interpolation. However, since different dimensions undergo different degrees of interpolation, the position embedding at the critical dimension will first become OOD. Thus, based on the scaled period of the critical dimension, the extrapolation upper bound $T_{\text{extra}}$ for NTK methods can be derived, as shown in Equation 5.
@@
Based on Equation 5, for an input length $t$, the rotary base scaling factor $\lambda$ should be set as shown in Equation 6 to ensure no OOD position embeddings occur. Notably, this adjustment coefficient exhibits a sup-linear, power-law increase with inference length (Liu et al., 2023b; 2025).
@@
It should be noted that while such interpolation could theoretically avoid extrapolation issues, it can only achieve 2$\times$ to 6$\times$ long-context extension during inference, as longer inputs lead to increased attention entropy, limiting further extrapolation (bloc97, 2023b; Han et al., 2023; Wang et al., 2024).
@@
Since the pre-trained position information of dimensions beyond the critical dimension limits the extrapolation capability of RoPE-based auto-regressive LLMs, if the rotary base is reduced, and each dimension can cover at least half or even a full period, the perplexity curve of auto-regressive LLMs will be flattened (Liu et al., 2023b). However, this does not imply real length extrapolation. Subsequent studies (Men et al., 2024; Hu et al., 2024) find that such LLMs can only perceive local information in downstream evaluations and fail to retrieve long-context dependencies.
Exposing LLM to periodic position information leads to downstream degeneration, manifesting a sliding-window effect, which reveals another aspect of RoPE-based extrapolation, the impact of monotonicity. Although higher dimensions do not observe complete position information, they provide a relatively complete monotonic interval, reflecting partial ordering in long-context scenarios. These dimensions exhibit larger activation values in long-context tasks (Jin et al., 2025), are more sensitive to modeling long-context dependencies (Liu et al., 2024a), and are better suited for capturing sequential relationships (Wei et al., 2025). Thus, solely optimizing for periodicity at the cost of losing monotonicity across all dimensions is wrong (Men et al., 2024; Liu et al., 2025).
@@ B.1 Setup
Since we have clarified the RoPE-based extrapolation in auto-regressive LLM, we now turn our focus to that in diffusion LLM and try to answer the three questions raised in Section 1. We conduct experiments on the existing diffusion LLM series, including LLaDA-8B (Nie et al., 2025), LLaDA-1.5 (Zhu et al., 2025), and Dream-v0 (Ye et al., 2025). By default, we set the number of sampling steps in diffusion LLM to 32 and keep the sampling strategy in the official code of LLaDA  [^1] https://github.com/ML-GSAI/LLaDA and Dream  [^2] https://github.com/HKUNLP/Dream. We use OpenCompass (Contributors, 2023) for validation. All experiments are performed with a fixed random seed of 2025, FP16 precision, and accelerated with FlashAttention2 (Dao, 2023).
B.2 More Results of Needle-In-A-Haystack
@@
Figure 8: NIAH Results of LLaDA-1.5 (Zhu et al., 2025) with different $\lambda$.
We report the NIAH results of LLaDA-8B-Instruct (Nie et al., 2025) and LLaDA (Zhu et al., 2025) with different sampling steps $s$ in Figure 11 and Figure 12 respectively, similar to Figure 3. In Figure 8, we report the NIAH results of pre-trained and NTK-scaled LLaDA-1.5 (Zhu et al., 2025). LLaDA-1.5 still supports 4k context length and has a local perception in direct length extrapolation. We use the same scaling factors as LLaDA and achieve similar length extrapolation effects consistent with the prediction of the scaling law of RoPE-based length extrapolation (Liu et al., 2023b).
In Figure 9 and Figure 10, we also report the NIAH performance of Dream-v0-7B-Base and Dream-v0-7B-Instruct. We find that both diffusion LLMs still have a local perception in direct extrapolation, but only support context length around 2k. We recalculate the scaling factor $\lambda$ for Dream-v0 in 4k, 8k, and 16k, respectively, based on Equation 1. Though the long-context capability of Dream-v0 Series is poor, we still find the NTK scaling works effectively in a 4k context length. Besides, since Dream-v0 Series is trained not from scratch but initialized from auto-regressive LLM, Qwen2.5-7B (Ye et al., 2025), we do not compare them with auto-regressive LLMs in task characteristics.
@@
(b) Dream-v0-7B-Base with $\lambda=5$
(c) Dream-v0-7B-Base with $\lambda=25$
(d) Dream-v0-7B-Base with $\lambda=126$
Figure 9: NIAH Results of Dream-v0-7B-Base (Ye et al., 2025) with different $\lambda$.
(a) Pretrained Dream-v0-7B-Instruct
(b) Dream-v0-7B-Instruct with $\lambda=5$
(c) Dream-v0-7B-Instruct with $\lambda=25$
(d) Dream-v0-7B-Instruct with $\lambda=126$
Figure 10: NIAH Results of Dream-v0-7B-Instruct (Ye et al., 2025) with different $\lambda$.
@@
Figure 11: NIAH Results of LLaDA-8B-Instruct (Nie et al., 2025) with different sampling steps.
@@
Figure 12: NIAH Results of LLaDA-1.5 (Zhu et al., 2025) with different sampling steps.

### v3

@@ 6 Related Work
Recently, Large Language Diffusion Models, or diffusion LLMs, have become a widely discussed topic in NLP research. After the theoretical simplification (sahoo2024simple; ou2024your; shi2024simplified) and empirical verification (he2022diffusionbert; gong2024scaling), researchers scale the size of diffusion LLMs to billions of parameters (nie2024scaling; nie2025large; dream2025) and demonstrate that diffusion LLMs can achieve comparable results with more promising performance in the reversal curse (berglund2023reversal). These immediately attract the attention of many more researchers. Significant research efforts have focused on adapting diffusion LLMs for multimodality, such as MMaDA (yang2025mmada), LLaDA-V (you2025llada), and LaViDa (li2025lavida), applying them to reasoning tasks, such as d1 (zhao2025d1), DCoLThuang2025reinforcing, and LLaDA-1.5 (zhu2025llada), and optimizing their efficiency (ma2025dkv; hu2025accelerating; wu2025fast), including dKV-Cache (ma2025dkv), Dimple (yu2025dimple), dLLM-Cache (liudllm), FreeCache (hu2025accelerating), Fast-dLLM (wu2025fast), and so on. However, there is still no discussion on the long-context capability of diffusion LLMs.
@@
Length extrapolation, or length generalization, or context extension, is an important issue for LLMs (press2021train; liu2025thus). The mainstream extrapolation research mainly focuses on adjusting position embedding, especially the widely used RoPE (su2021roformer). For example, Linear PI (chen2023extending) first achieves LLMs’ length extrapolation by scaling position indices to the pre-training range with little fine-tuning. The NTK method (fixedNTK; dynamicNTK; peng2023yarn) then scales the rotary base in RoPE (su2021roformer) to achieve plug-and-play length extrapolation. Subsequently, amplifying the rotary base and training on longer lengths has become the dominant approach for length extrapolation (roziere2023code; xiong2023effective; liu2023scaling; ding2024longrope). In addition, ReRoPE (rerope), ReAttention (liu2024reattention), and DCA (an2024training; an2024does) also achieve plug-and-play extrapolation by limiting the relative position. In this paper, we still focus on the length extrapolation via NTK scaling (fixedNTK; liu2023scaling) in the inference stage, and try to reveal and explain the similarities and differences in length extrapolation between diffusion-based and auto-regressive LLM.
@@ Appendix A Preliminary: RoPE Extrapolation in Auto-regressive LLM
Rotary Position Embedding (RoPE) (su2021roformer) employs trigonometric functions to encode absolute positions in Q state $\bm{q}_{t}=\left[q_{t}^{(0)},\cdots,q_{t}^{(d-1)}\right]$ and K state $\bm{k}_{s}=\left[k_{s}^{(0)},\cdots,k_{s}^{(d-1)}\right]$. By leveraging the properties of rotation matrices, RoPE encodes relative position in the attention matrix $\bm{A}$, as shown in Equation 2 and demonstrates superior performance, thus being widely adopted by many auto-regressive LLMs (Sun2024MOSS; dubey2024llama; qwen2).
$\begin{aligned} \bm{A}_{t,s}&=\left(\bm{q}_{t}\bm{R}_{t}\right)\left(\bm{k}_{s}\bm{R}_{s}\right)^{\top}=\bm{q}_{t}\bm{R}_{t-s}\bm{k}_{s}^{\top}\\ &=\sum\limits_{n=0}^{d/2-1}{\begin{bmatrix}q_{t}^{(2n)}\\ q_{t}^{(2n+1)}\end{bmatrix}^{\top}\begin{bmatrix}\cos{\theta_{n}(t-s)}&-\sin{\theta_{n}(t-s)}\\ \sin{\theta_{n}(t-s)}&\cos{\theta_{n}(t-s)}\end{bmatrix}\begin{bmatrix}k_{s}^{(2n)}\\ k_{s}^{(2n+1)}\end{bmatrix}}\\ &=\sum\limits_{n=0}^{d/2-1}{\begin{gathered}\left(q_{t}^{(2n)}k_{s}^{(2n)}+q_{t}^{(2n+1)}k_{s}^{(2n+1)}\right)\cos{\theta_{n}(t-s)}\\ -\left(q_{t}^{(2n)}k_{s}^{(2n+1)}-q_{t}^{(2n+1)}k_{s}^{(2n)}\right)\sin{\theta_{n}(t-s)}\end{gathered}}\end{aligned}\text{.}$ (2)
However, RoPE still faces the length extrapolation issue (press2021train). When RoPE-based auto-regressive LLMs are tested beyond the pre-trained context length, the perplexity rises significantly, and downstream performance drops sharply. The underlying causes and corresponding solutions can be attributed to two key properties of trigonometric functions: periodicity and monotonicity.
@@
Therefore, dimensions beyond $d_{\text{extra}}$ will encounter OOD position embedding when processing longer inputs and larger position indices in inference, leading to extrapolation issues (liu2023scaling).
To enable LLM to handle unseen position indices, NTK methods (fixedNTK; xiong2023effective) scale the rotary base by a factor $\lambda$, reducing the rotary angle to achieve position interpolation. However, since different dimensions undergo different degrees of interpolation, the position embedding at the critical dimension will first become OOD. Thus, based on the scaled period of the critical dimension, the extrapolation upper bound $T_{\text{extra}}$ for NTK methods can be derived, as shown in Equation 5.
@@
Based on Equation 5, for an input length $t$, the rotary base scaling factor $\lambda$ should be set as shown in Equation 6 to ensure no OOD position embeddings occur. Notably, this adjustment coefficient exhibits a sup-linear, power-law increase with inference length (liu2023scaling; liu2025thus).
@@
[…]
@@
Since the pre-trained position information of dimensions beyond the critical dimension limits the extrapolation capability of RoPE-based auto-regressive LLMs, if the rotary base is reduced, and each dimension can cover at least half or even a full period, the perplexity curve of auto-regressive LLMs will be flattened (liu2023scaling). However, this does not imply real length extrapolation. Subsequent studies (men2024base; hu2024can) find that such LLMs can only perceive local information in downstream evaluations and fail to retrieve long-context dependencies.
Exposing LLM to periodic position information leads to downstream degeneration, manifesting a sliding-window effect, which reveals another aspect of RoPE-based extrapolation, the impact of monotonicity. Although higher dimensions do not observe complete position information, they provide a relatively complete monotonic interval, reflecting partial ordering in long-context scenarios. These dimensions exhibit larger activation values in long-context tasks (jin2025massive), are more sensitive to modeling long-context dependencies (liu2024beyond), and are better suited for capturing sequential relationships (wei2025videorope). Thus, solely optimizing for periodicity at the cost of losing monotonicity across all dimensions is wrong (men2024base; liu2025thus).
@@ B.1 Setup
Since we have clarified the RoPE-based extrapolation in auto-regressive LLM, we now turn our focus to that in diffusion LLM and try to answer the three questions raised in Section 1. We conduct experiments on the existing diffusion LLM series, including LLaDA-8B (nie2025large), LLaDA-1.5 (zhu2025llada), and Dream-v0 (dream2025). By default, we set the number of sampling steps in diffusion LLM to 32 and keep the sampling strategy in the official code of LLaDA  [^1] https://github.com/ML-GSAI/LLaDA and Dream  [^2] https://github.com/HKUNLP/Dream. We use OpenCompass (2023opencompass) for validation. All experiments are performed with a fixed random seed of 2025, FP16 precision for LLaDA Series and BF16 for Dream Series, and accelerated with FlashAttention2 (dao2023flashattention).
B.2 More Experiment Results
@@
Figure 8: NIAH Results of LLaDA-1.5 (zhu2025llada) with different $\lambda$.
We report the NIAH results of LLaDA-8B-Instruct (nie2025large) and LLaDA (zhu2025llada) with different sampling steps $s$ in Figure 10 and Figure 11 respectively, similar to Figure 3. In Figure 8, we report the NIAH results of pre-trained and NTK-scaled LLaDA-1.5 (zhu2025llada). LLaDA-1.5 still supports 4k context length and has a local perception in direct length extrapolation. We use the same scaling factors as LLaDA and achieve similar length extrapolation effects consistent with the prediction of the scaling law of RoPE-based length extrapolation (liu2023scaling).
In Figure 9, we also report the NIAH performance of Dream-v0-7B-Base and Dream-v0-7B-Instruct. We find that both diffusion LLMs still have a local perception in direct extrapolation. Since Dream-v0 Series is trained not from scratch but initialized from an auto-regressive LLM, Qwen2.5-7B (dream2025), it is hard to define the exact context length of pre-training. Therefore, we only compare the long-context performance of pretrained models to verify the task characteristics as shown in Table 3 and Table 4. Dream Series achieves a comparable and even better performance in LongBench (bai2023longbench), and demonstrates strength in synthetic QA and weakness in aggregation tasks in RULER (hsieh2024ruler) within a 4k context.
@@
(b) Pretrained Dream-v0-7B-Instruct
Figure 9: NIAH Results of Dream-v0-7B Series (dream2025).
4k 8k
SD MD Sum ICL Syn Code Avg SD MD Sum ICL Syn Code Avg
Dream-v0-7B-Base 22.8 20.7 34.1 36.3 64.4 68.3 38.8 24.0 21.6 35.9 40.3 73.9 68.9 41.7
Dream-v0-7B-Instruct 23.8 21.7 35.1 37.3 65.4 69.3 39.8 25.0 22.6 36.9 41.3 74.9 69.9 42.7
Table 3: Results of Dream-v0-7B (dream2025) and LLaMA3-8B (meta2024llama) on LongBench (bai2023longbench) under 4k and 8k context length.
4k 8k 16k
NIAH AGG QA Avg NIAH AGG QA Avg NIAH AGG QA Avg
Dream-v0-7B-Base 90.6 68.3 78.5 83.6 68.1 49.7 61.5 62.9 24.3 26.5 45.0 28.0
Dream-v0-7B-Instruct 95.6 74.6 76.5 87.8 72.3 55.8 58.0 66.3 31.0 29.5 46.0 33.0
LLaMA3-8B-Base 99.8 98.1 67.5 94.4 99.6 93.5 63.0 92.5 0.0 0.0 0.0 0.0
Table 4: Results of Dream-v0-7B (dream2025) and LLaMA3-8B (meta2024llama) on RULER (hsieh2024ruler) under 4k, 8k and 16k context length.
@@
Figure 10: NIAH Results of LLaDA-8B-Instruct (nie2025large) with different sampling steps.
@@
Figure 11: NIAH Results of LLaDA-1.5 (zhu2025llada) with different sampling steps.